• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Keppel

    Log entry time 16:05:42 on November 3,2006

    Entry number 188811

    keyword=Shift Summary

    The shift started smoothly running with the end of run 1925. We continued until ~8:50 AM when MCC had an ISD problem.

    The end run failed for 1926, and we had CODA problems. Fixed by a ROC reboot and a CODA kill and restart, and we were back up running at 1927 by 9:05 AM.

    At about 10:30 AM, we stopped beam with run 1928 as the target was having large temperature and pressure flucuations. It turns out this was due to the cryo group readjusting their liquid levels, and so we requested beam again at about 10:40 AM.

    After a few minutes of accelerator problems we started run 1929. This was a bust, as accelerator kept having problems. This went on until about 11:10 AM.

    We took data until a brief stop for Hall C to take out the half wave plate at 11:55 AM. There was an accelerator problem at the same time.

    We started again with run 1932 at 12:03 PM.

    Run 1933 stopped ~1:10 PM for our beam energy measurement.
    Target out, raster off.

    The beam energy measurement finished at 3:25 PM, and we began a Moller measurement....